OpenStudy (kiiraa_x3):

If a parabola's focus is at (−2, 5) and the directrix is at y = −1, what is the equation representing this parabola? x2 + 4x − 12y + 67 = 0 x2 − 28 = 0 x2 − 4x + 12y − 28 = 0 x2 + 4x − 12y + 28 = 0

OpenStudy (kiiraa_x3):

@Michele_Laino @pooja195 @Hero @triciaal

OpenStudy (anonymous):

you can use the definition of a parabola

OpenStudy (anonymous):

distance ( -2,5) ( x,y) = distance ( ( x,y) , ( x, -1) )

OpenStudy (anonymous):



OpenStudy (anonymous):

y = 1/12 ( (x+2)^2 + 24) 12y = (x+2)^2 + 24 12y = x^2 + 4x + 4 + 24 x^2 + 4x - 12y +28= 0

OpenStudy (anonymous):

ded. no response